OpenStudy (anonymous):

find the exact value of cos (2 theta - pi over 2)

OpenStudy (akashdeepdeb):

One rule we need to know is: \(cos (-\theta) = cos\theta \) You are given: \[cos (2\theta - \frac{\pi}{2})\] Take the minus sign common: You'll get; \[cos (-(-2\theta + \frac{\pi}{2})) = cos (\frac{\pi}{2}-2\theta)\] cos (90-a) = sin a Thus, \[cos (\frac{\pi}{2}-2\theta) = sin 2\theta\] Understood? :)